I would like for you to help me with this problem:
The Everton college store paid $1768 for an order of 47 calculators. the store paid $8 for each scientific calculator. The others, all graphing calculators cost the store $56 each. How many of each type of calculator was ordered?

Answer by Earlsdon(6103)

Let x = the number of graphing calculators at $56 each, then (47-x) = the number of scientific calclators at $8 each.
The cost of each group of calculators can be expressed as:
x graphing calculators at $56 each = ($56)*x.
(47-x) scientific calculators at $8 each = ($8)*(47-x).
The sum of these two amounts is given as $1,768.00 so we have enough information to form an equation in x.
(56)(x) + (8)*(47-x) = 1769 Simplify and solve for x.
56x+376-8x = 1768 Combine the x-terms on the left.
48x+376 = 1768 Subtract 376 from both sides.
48x = 1393 Divide both sides by 48.
x = 29 and...
(47-x) = 47-29 = 18.
29 graphing calculators and 18 scientific calculators were ordered.